# Neural Context Flows for Meta-Learning of Dynamical Systems

### v5

@@ B.4 Lotka-Volterra (LV)
With dynamics that are of continued interest to many fields (epidemiology Venturino (1994), economy Wu et al. (2012), etc.), the Lotka-Volterra (LV) ODE models the evolution of the concentration of preys $x$ and predators $y$ in a closed ecosystem. The behavior of the system is controlled by the prey’s natural growth rate $\alpha$, the predation rate $\beta$, the predator’s increase rate upon consuming prey $\delta$, and the predator’s natural death rate $\gamma$
@@ B.6 Sel’kov Model (SM)
Figure 18: Per-environment In-Domain and adaptatation MSEs for the SM problem. In blue are the meta-training environments, in red the meta-testing ones. Central environments in the attractor (E) are better resolved than the others.
@@ B.7 Brusselator (BT)
  • $A$ and $B$ are constants representing the rate parameters of the chemical reactions, both varying across environments.
Using an RK4 adaptive time-step solver, the system is simulated up to $T=10$ (excluded), with the step reported every time step $\Delta t=0.5$. The initial condition for each trajectory for all environments is sampled and broadcasted as follows:
@@
As observed with the metrics in Table 2, the BT dataset is one of the most challenging to learn on. Indeed, the non-Meta-Learning baselines OFA and OPE equally struggle on it. To show this, we vary the number of trajectories in each trainnig environment between 1 and 8. Then, we plot in Fig. 19 the sample efficiency of our NCF approach against the non-Meta-Learning baselines. Like with the SP problem (cf. Fig. 12), we observe good NCF performance when data is scarce, underlining the suitability of our approach for few-shot learning.
Figure 19: In-Domain MSEs on the BT problem comparing the sample eficiency of NCF-$\mathit{t}_{2}$ against OPE and OFA. NCF is effective in low-data regimes, and OPE closes the gap as data increases.
@@ C.1 Limiting the context pool size
The context pool $\mathrm{P}$ from which environments $j$ are randomly sampled contributes significantly to the computational and memory complexity of the algorithm. As evidenced in Algorithms 3 and 1, the computation of the loss and hence its gradient can be parallelized across the contexts vectors $\xi^{j}$. With the goal of assessing the associated computational burden, we vary the size of the pool size for the LV problem from 1 to 9, reporting the In-Domain and OoD metrics, and computational time in LABEL:fig:contextpoolsize.
We dig further into the influence of the context pool size. While its ideal value might be difficult to conclude based on LABEL:fig:contextpoolsize, LABEL:fig:contextpoolsize_T1 investigates NCF-$\mathit{t}_{1}$ to highlight how important any value bigger than 1 is (with $p=6$ proving to be exceptionally adequate). Indeed, 0* indicates no context pool was used, and the vector field is evaluated without Taylor expansion. This ablation results in a significant drop in MSE of about one order of magnitude. While computational meta-training time scales linearly with the pool size (see LABEL:fig:poolsizetimes), the various losses do not. Overall, these results suggest that the context pool size $p$ largely remains a hyperparameter that should be tuned for maximum balance of accuracy and training time.
Our motivations for the above conclusion lies in that Taylor expansion is the key to information flowing from one environment to another, along with the concept of “task relatedness” or “context proximity”. Indeed, depending on the pool-filling strategy used, some environments that are prohibitively far from one another may be required to interact in context space. As a result, the Taylor approximation incurs a non-reducible residual error term. To counter this effect, one could restrict the pool size to limit the impact of those far-apart environments since they will be less often sampled from the pool as the training progresses.
@@ C.3 Variation of the Pooling Strategy
Although all strategies capture the underlying structure of the LV physical system (see Fig. 21), Table 6 indicates that RA and NF have the best performance, contrasting with high metrics and uncertainties displayed by SF. The training and validation dynamics in Fig. 22 provide clearer insight into this discrepancy, highlighting an important spread of validation values for SF. Interestingly, Fig. 22 shows that RA takes much longer to converge during training, despite ultimately producing the best OoD performance. This slow convergence is because very early on, the model is still forced to find commonalities between all pairs of environment, despite some pairs being less related than others. All in all, NF is the most balanced strategy on the LV problem since it converges fast, shows signs of increased stability, and produces excellent results InD and OoD.
Table 6: Comparison of pooling strategies on the LV problem. The reported mean and standard deviation training and sequential adaptation times are expressed in minutes. All strategies are trained for the same number of epochs across 3 separate runs with identical hyperparameters.
@@ C.4 Ablation of the 3-Networks Architecture
This study is of particular significance since, if in addition to foregoing the 3-networks architecture, we eliminated the Taylor expansion step, and NCF would turn into the data-controlled neural ODE Massaroli et al. (2020). So we run both dynamics forecasting problems without the 3-networks architecture, and we report the training MSE in Fig. 23. In-domain and adaptation MSEs for both LV and SP problems are reported in Table 7.
@@ Appendix D Experimental Details
Our main workstation for training and adaptation was fitted with an Nvidia GeForce RTX 4080 graphics card which was used for the SP, LV, SM, and NS problems. Additionally, we used an RTX 3090 GPU of the same generation for the GO problem, and an NVIDIA A100 Tensor Core GPU for the GS problem as its CNN-based architecture was more memory-intensive. In addition to the training hardware and deep learning frameworks (JAX for NCFs, and PyTorch for CAVIA and CoDA) that made apple-to-apple comparison challenging, our network architectures varied slightly across methods, as the subsections below highlight. This is part of the reason we decided to focus on total parameter count as the great equalizer.
@@ D.1 NCFs
The MLPs used as our context, state and main networks (as depicted in Fig. 2b) typically had a uniform width. When that was not the case, we use the notation $[h_{\text{in}}\rightarrow h_{1}\rightarrow...\rightarrow h_{d-1}\rightarrow h_{\text{out}}]$ to summarize the number of units in each layer of an MLP of depth $d$ for brevity. We used a similar notation to summarize the number of channels in a Convolutional Neural Network (CNN). For instance, on the SP problem, the context network was the MLP $[256\rightarrow 64\rightarrow 64\rightarrow 64]$ (which indicates a context size of $d_{\xi}=256$), the state network was $[2\rightarrow 64\rightarrow 64\rightarrow 64]$, and the main network was $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$, all with Swish activations Ramachandran et al. (2017). For the OFA and OPE experiments that do not require contextual information (cf. Tables 1 and 4), we delete the state and context networks then increase the hidden units of the main network to $156$ to match the NCF parameter count. Below, we follow the same convention to detail the networks used for other problems in this paper’s main comparison (see Table 2).
@@
LV (context) $[1024\rightarrow 256\rightarrow 256\rightarrow 64]$; (state) $[2\rightarrow 64\rightarrow 64\rightarrow 64]$; (main) $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GO (context) $[256\rightarrow 64\rightarrow 64\rightarrow 122]$; (state) $[7\rightarrow 122\rightarrow 122\rightarrow 122]$; (main) $[244\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 7]$SM (context, state, main) Exact same architectures as with SP above • BT (context) linear layer with 256 input and 128 output features followed by an appropriate reshaping then a 2D convolution with 8 channels and circular padding to maintain image size, a kernel of size $3\times 3$, and Swish activation [^12] Except for the number of channels, these were default convolution settings we used throughout. The circular padding enforced the periodic boundary condition observed in the trajectory data.; (state) appropriate reshaping before 2D convolution with 8 channels; (main) CNN $[16\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GS Same network architectures as for the previous BT problem, with the exception that the linear layer in the context network outputted 2048 features • NS (context) A single linear layer with 202 input and 1024 output features reshaped and concatenated to the state and the grid coordinates before processing; (state) No state network was used for this problem; (main) 2-dimensional Fourier Neural Operator (FNO) Li et al. (2020) with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator was a $1\times 1$ convolution with 4 input and 10 output channels respectively – functionally similar to a fully connected layer with weight sharing. The projection operator had two such layers, with 10 inputs, 1 output, and 16 hidden channels in-between.
@@
As for the propagation of gradients from the loss function to the learnable parameters in the right-hand-side of the neural ODEs (5), we opted to differentiate through our solvers, rather than numerically solving for adjoint states Chen et al. (2018). Since our main solvers were the less demanding RK4 and Dopri5 solvers, we didn’t incur the heavy memory cost Differentiable Programming methods are typically known for Nzoyem et al. (2023); Kidger et al. (2020).
Specifically, we used the following integration schemes: Dopri5 Wanner & Hairer (1996) for GO, SM, GS, and BT, with relative and absolute tolerances of $10^{-3}$ and $10^{-6}$ respectively; RK4 for LV with $\Delta t=0.1$; Explicit Euler for NS with $\Delta t=1$. Across all NCF variants, we kept the context pool size under 4 to reduce computational workload ($p=2$ for LV and SM, and $p=4$ for LV and GO, and $p=3$ for all PDE problems). We used the Adam optimizers Diederik (2014) for both model weights and contexts on ODE problems, and Adabelief Zhuang et al. (2020) for PDE problems. Their initial learning rates were as follows: $3\times 10^{-4}$ for LV and NS, $10^{-3}$ for GO, BT, and GS, $10^{-4}$ for SM. That learning rate was kept constant throughout the various trainings, except for BT, GS, and NS where it was multiplied by a factor (0.1, 0.5, and 0.1 respectively) after a third of the total number of training steps, and again by the same factor at two-thirds completion. The same initial learning rates were used during adaptation, with the number of iterations typically set to 1500.
@@ Appendix E Example Implementation of NCFs
The vector field takes center-stage when modeling using ODEs. It is critical to use JVPs, since we never want to materialize the Jacobian (the context size can be prohibitively large). The Jacobian-Vector Product primitive from Equinox Kidger & Garcia (2021) filter_jvp as illustrated below, coupled with jit-compilation, enabled fast runtimes for all problems.
@@
Eq. 10 provides a structured summation layout particularly suited for a function transformation like JAX’s vmap. During implementation, that loss functions can be defined in two stages, the innermost summation term along with Eq. 9 making up the first stage; while the two outermost summations of equation 10 make up the second stage. Once the first stage is complete, the second is relatively easy to implement. We provide a vectorized implementation of the first stage below.

### v6

@@ B.4 Lotka-Volterra (LV)
With dynamics that are of continued interest to many fields (epidemiology Venturino (1994), economy Wu et al. (2012), etc.), the Lotka-Volterra (LV) ODE models the evolution of the concentration of prey $x$ and predators $y$ in a closed ecosystem. The behavior of the system is controlled by the prey’s natural growth rate $\alpha$, the predation rate $\beta$, the predator’s increase rate upon consuming prey $\delta$, and the predator’s natural death rate $\gamma$
@@
(a) Results for the Large adaptation of the LV problem to a grid, showcasing the MAPE and training MSE. Compared to Fig. 3, the colorbar range in (a) is shrunk to focus on the MAPEs between 0 and 8%.
@@ B.6 Sel’kov Model (SM)
Figure 18: Per-environment In-Domain and adaptation MSEs for the SM problem. In blue are the meta-training environments, in red the meta-testing ones. Central environments in the attractor (E) are better resolved than the others.
@@ B.7 Brusselator (BT)
  • AA and $B$ are constants representing the rate parameters of the chemical reactions, both varying across environments.
Using an RK4 adaptive time-step solver, the system is simulated up to $T=10$ (excluded), with the step reported every time step $\Delta t=0.5$. The initial condition for each trajectory for all environments is sampled and broadcast as follows:
@@
As observed with the metrics in Table 2, the BT dataset is one of the most challenging to learn on. Indeed, the non-Meta-Learning baselines OFA and OPE equally struggle with it. To show this, we vary the number of trajectories in each training environment between 1 and 8. Then, we plot in Fig. 19 the sample efficiency of our NCF approach against the non-Meta-Learning baselines. Like with the SP problem (cf. Fig. 12), we observe good NCF performance when data is scarce, underlining the suitability of our approach for few-shot learning.
Figure 19: In-Domain MSEs on the BT problem comparing the sample efficiency of NCF-$\mathit{t}_{2}$ against OPE and OFA. NCF is effective in low-data regimes, and OPE closes the gap as data increases.
@@ C.1 Limiting the context pool size
The context pool $\mathrm{P}$ from which environments $j$ are randomly sampled contributes significantly to the computational and memory complexity of the algorithm. As evidenced in Algorithms 3 and 1, the computation of the loss and hence its gradient can be parallelized across the context vectors $\xi^{j}$. With the goal of assessing the associated computational burden, we vary the size of the pool size for the LV problem from 1 to 9, reporting the In-Domain and OoD metrics, and computational time in Fig. 21(a).
(a) (a) In-domain and OoD MSEs, and (b) training and adaptation times when varying the context pool size on the LV problem with NCF-$\mathit{t}_{2}$.
(b) Average MSEs for the ablation of the context pool size for NCF-$\mathit{t}_{1}$, with evaluation carried over many seeds, hence the vertical bars indicating the standard deviations.
We dig further into the influence of the context pool size. While its ideal value might be difficult to conclude based on Fig. 21(a), Fig. 21(b) investigates NCF-$\mathit{t}_{1}$ to highlight how important any value bigger than 1 is (with $p=6$ proving to be exceptionally adequate). Indeed, 0* indicates no context pool was used, and the vector field is evaluated without Taylor expansion. This ablation results in a significant drop in MSE of about one order of magnitude. While computational meta-training time scales linearly with the pool size (see LABEL:fig:poolsizetimes), the various losses do not. Overall, these results suggest that the context pool size $p$ largely remains a hyperparameter that should be tuned for maximum balance of accuracy and training time.
Our motivations for the above conclusion lie in the fact that Taylor expansion is the key to information flowing from one environment to another, along with the concept of “task relatedness” or “context proximity”. Indeed, depending on the pool-filling strategy used, some environments that are prohibitively far from one another may be required to interact in context space. As a result, the Taylor approximation incurs a non-reducible residual error term. To counter this effect, one could restrict the pool size to limit the impact of those far-apart environments since they will be less often sampled from the pool as the training progresses.
@@ C.2 Restriction of the context size
(c) (a) InD and OoD evaluation MSEs for the LV problem as the context size is increased. (b) Loss landscape plotted in context vectors space for $d_{\xi}=2$. The clustering of the training contexts mirrors the physical parameter pairs $(\beta,\delta)$ from Fig. 3.
@@ C.3 Variation of the Pooling Strategy
Although all strategies capture the underlying structure of the LV physical system (see Fig. 21), Table 6 indicates that RA and NF have the best performance, contrasting with high metrics and uncertainties displayed by SF. The training and validation dynamics in Fig. 22 provide clearer insight into this discrepancy, highlighting an important spread of validation values for SF. Interestingly, Fig. 22 shows that RA takes much longer to converge during training, despite ultimately producing the best OoD performance. This slow convergence is because very early on, the model is still forced to find commonalities between all pairs of environments, despite some pairs being less related than others. All in all, NF is the most balanced strategy on the LV problem since it converges fast, shows signs of increased stability, and produces excellent results InD and OoD.
Table 6: Comparison of pooling strategies on the LV problem. The reported mean and standard deviation of training and (sequential) adaptation times are expressed in minutes. All strategies are trained for the same number of epochs across 3 separate runs with identical hyperparameters.
@@ C.4 Ablation of the 3-Networks Architecture
This study is of particular significance since, if in addition to foregoing the 3-networks architecture, we eliminated the Taylor expansion step, NCF would turn into the data-controlled neural ODE Massaroli et al. (2020). So we run both dynamics forecasting problems without the 3-networks architecture, and we report the training MSE in Fig. 23. In-domain and adaptation MSEs for both LV and SP problems are reported in Table 7.
@@ Appendix D Experimental Details
Our main workstation for training and adaptation was fitted with an Nvidia GeForce RTX 4080 graphics card which was used for the SP, LV, SM, and NS problems. Additionally, we used an RTX 3090 GPU of the same generation for the GO problem, and an NVIDIA A100 Tensor Core GPU for the GS problem, as its CNN-based architecture was more memory-intensive. In addition to the training hardware and deep learning frameworks (JAX for NCFs, and PyTorch for CAVIA and CoDA) that made apple-to-apple comparison challenging, our network architectures varied slightly across methods, as the subsections below highlight. This is part of the reason we decided to focus on total parameter count as the great equalizer.
@@ D.1 NCFs
The MLPs used as our context, state and main networks (as depicted in Fig. 2b) typically had a uniform width. When that was not the case, we use the notation $h_{\text{in}}\rightarrow h_{1}\rightarrow...\rightarrow h_{d-1}\rightarrow h_{\text{out}}]$ to summarize the number of units in each layer of an MLP of depth $d$ for brevity. We used a similar notation to summarize the number of channels in a Convolutional Neural Network (CNN). For instance, on the SP problem, the context network was the MLP $256\rightarrow 64\rightarrow 64\rightarrow 64]$ (which indicates a context size of $d_{\xi}=256$), the state network was $2\rightarrow 64\rightarrow 64\rightarrow 64]$, and the main network was $128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$, all with Swish activations Ramachandran et al. (2017). For the OFA and OPE experiments that do not require contextual information (cf. Tables 1 and 4), we delete the state and context networks then increase the hidden units of the main network to $156$ to match the NCF parameter count. Below, we follow the same convention to detail the networks used for other problems in this paper’s main comparison (see Table 2).
@@
LV (context) $1024\rightarrow 256\rightarrow 256\rightarrow 64]$; (state) $2\rightarrow 64\rightarrow 64\rightarrow 64]$; (main) $128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GO (context) $256\rightarrow 64\rightarrow 64\rightarrow 122]$; (state) $7\rightarrow 122\rightarrow 122\rightarrow 122]$; (main) $244\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 7]$SM (context, state, main) Exact same architectures as with SP above • BT (context) linear layer with 256 input and 128 output features followed by an appropriate reshaping then a 2D convolution with 8 channels and circular padding to maintain image size, a kernel of size $3\times 3$, and Swish activation [^12] Except for the number of channels, these were default convolution settings we used throughout. The circular padding enforced the periodic boundary condition observed in the trajectory data.; (state) appropriate reshaping before 2D convolution with 8 channels; (main) CNN $16\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GS Same network architectures as for the previous BT problem, with the exception that the linear layer in the context network outputted 2048 features • NS (context) A single linear layer with 202 input and 1024 output features reshaped and concatenated to the state and the grid coordinates before processing; (state) No state network was used for this problem; (main) 2-dimensional Fourier Neural Operator (FNO) Li et al. (2020) with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator was a $1\times 1$ convolution with 4 input and 10 output channels respectively – functionally similar to a fully connected layer with weight sharing. The projection operator had two such layers, with 10 inputs, 1 output, and 16 hidden channels in-between.
@@
As for the propagation of gradients from the loss function to the learnable parameters in the right-hand side of the neural ODEs (5), we opted to differentiate through our solvers, rather than numerically solving for adjoint states Chen et al. (2018). Since our main solvers were the less demanding RK4 and Dopri5 solvers, we didn’t incur the heavy memory cost that Differentiable Programming methods are typically known for Nzoyem et al. (2023); Kidger et al. (2020).
Specifically, we used the following integration schemes: Dopri5 Wanner & Hairer (1996) for GO, SM, GS, and BT, with relative and absolute tolerances of $10^{-3}$ and $10^{-6}$ respectively; RK4 for LV with $\Delta t=0.1$; Explicit Euler for NS with $\Delta t=1$. Across all NCF variants, we kept the context pool size under 4 to reduce computational workload ($p=2$ for LV and SM, and $p=4$ for LV and GO, and $p=3$ for all PDE problems). We used the Adam optimizers Diederik (2014) for both model weights and contexts on ODE problems, and Adabelief Zhuang et al. (2020) for PDE problems. Their initial learning rates were as follows: $3\times 10^{-4}$ for LV and NS, $10^{-3}$ for GO, BT, and GS, $10^{-4}$ for SM. That learning rate was kept constant throughout the various trainings, except for BT, GS, and NS where it was multiplied by a factor (0.1, 0.5, and 0.1, respectively) after a third of the total number of training steps, and again by the same factor at two-thirds completion. The same initial learning rates were used during adaptation, with the number of iterations typically set to 1500.
@@ Appendix E Example Implementation of NCFs
The vector field takes center stage when modeling using ODEs. It is critical to use JVPs, since we never want to materialize the Jacobian (the context size can be prohibitively large). The Jacobian-Vector Product primitive from Equinox Kidger & Garcia (2021) filter_jvp as illustrated below, coupled with jit-compilation, enabled fast runtimes for all problems.
@@
Eq. 10 provides a structured summation layout particularly suited for a function transformation like JAX’s vmap. During implementation, the loss functions can be defined in two stages, the innermost summation term along with Eq. 9 making up the first stage; while the two outermost summations of equation 10 make up the second stage. Once the first stage is complete, the second is relatively easy to implement. We provide a vectorized implementation of the first stage below.